A pharmacist is to prepare 18 milliliters of special eye drops for a glaucoma patient. The eye drop solution must have a 5% active ingredient, but the pharmacist only has 13% solution and 3% solution in stock.
a) How much of the 13% solution should be used in the prescription?
b) How much of the 3% solution should be used in the prescription?

Answer by edjones(8007)

Eye drops are dispensed in milliliters NOT millimeters. I corrected the question for you. Abbreviation is ml.
.13 * x = .13x
.03*(18-x)=.54-.03x
.05 * 18 = .9
.
.13x-.03x+.54=.9
.1x+.54=.9
.1x=.36
x=3.6 ml.
18-x=14.4 ml.
.
Ed